### v1

# LLMInit: A Free Lunch from Large Language Models for Selective Initialization of Recommendation

Abstract.
Collaborative filtering models, particularly graph-based approaches, have demonstrated strong performance in capturing user-item interactions for recommendation systems. However, they continue to struggle in cold-start and data-sparse scenarios. The emergence of large language models (LLMs) like GPT and LLaMA presents new possibilities for enhancing recommendation performance, especially in cold-start settings. Despite their promise, LLMs pose challenges related to scalability and efficiency due to their high computational demands and limited ability to model complex user-item relationships effectively. In this work, we introduce a novel perspective on leveraging LLMs for CF model initialization. Through experiments, we uncover an embedding collapse issue when scaling CF models to larger embedding dimensions. To effectively harness large-scale LLM embeddings, we propose innovative selective initialization strategies utilizing random, uniform, and variance-based index sampling. Our comprehensive evaluation on multiple real-world datasets demonstrates significant performance gains across various CF models while maintaining a lower computational cost compared to existing LLM-based recommendation approaches.
Collaborative Filtering, Recommendation, Large Language Models
doi: XXXXXXX.XXXXXXXccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
1. Introduction
Recommender systems (RecSys), particularly collaborative filtering (CF), play a pivotal role in many online platforms and web applications  (Koren et al., 2021; Fan et al., 2019; Zou et al., 2024), aiding users in navigating vast information by offering personalized suggestions. Among the various techniques, collaborative filtering (Wang et al., 2019; He et al., 2020; Xu et al., 2023; Zhang et al., 2024a; Xu et al., 2025) stands out for its effectiveness in predicting user preferences by analyzing past user-item interactions. A key principle among these models is learning randomly initialized user/item representations with carefully designed loss functions to capture each user’s preference (Rendle et al., 2012). However, such CF approaches heavily rely on the observed user-item interactions to optimize the CF embeddings from random initialization and perform poorly in real-world applications where interaction data is sparse and users/items are cold-start.
More recently, the rapid advancement of large language models (LLMs) such as GPT-4  (Achiam et al., 2023) and LLaMa  (Touvron et al., 2023), which have exhibited remarkable proficiency in understanding and processing textual information, have piqued the interest of researchers seeking to explore their potential in improving recommender systems beyond traditional approaches  (Geng et al., 2022; Yuan et al., 2023; Li et al., 2023). One of the most promising directions in this exploration involves adapting LLMs as recommender systems through prompt engineering and tailored instruction tuning. Such integration of LLMs enables the RecSys to better understand complex user preferences, and interpret contextual information, thereby surpassing the traditional models in cold-start settings.
Figure 1. An investigation of the embedding collapse issue of CF model LightGCN (He et al., 2020) in two Amazon datasets. Increasing embedding size with exponentially growing parameters finally leads to performance degradation.
Despite the inspiring progress made in LLMs for recommendations, their adoption in online deployments faces significant challenges related to efficiency and scalability. These issues stem from the inherently time-consuming and computationally intensive nature of LLMs. Such problems can be further exaggerated in real-world applications of large-scale users and items due to the enlarged vocabulary tokens and in-context input. Moreover, LLMs fall short of complex CF interaction understanding owing to the next-token prediction pipeline and the constraints of a maximum token limit (e.g., 4096 tokens in LLaMA) to incorporate the whole CF interactions. This limitation hampers their ability to capture and effectively model global user-item dependencies as in CF models, leading to poor performance in the full-ranking (Hou et al., 2024) and warm-start settings (Bao et al., 2023). Therefore, we raise a critical research question:
How can we harness the power of LLMs for recommendation systems in an effective and scalable manner?
To answer the above question, we propose to utilize LLMs for CF-based RecSys from a new perspective by taking pretrained large language models as a free lunch for the embedding initialization of CF models. One trivial way is to directly utilize the LLMs to generate embeddings with substantial textual information. However, storing the large-scale embeddings from LLMs for each user and item is not efficient nor scalable, especially when the platform scales to millions of users and items. Furthermore, as illustrated in Figure 1, unlike the scaling laws observed in NLP models (Kaplan et al., 2020), the performance of the recommendation model LightGCN (He et al., 2020) deteriorates even when the numbers of parameters scale up exponentially. This behavior resembles the embedding collapse issue (Guo et al., [n. d.]) observed in click-through rate prediction models (Guo et al., 2017; Wang et al., 2021). These findings suggest that significantly smaller embedding sizes (e.g., 128) are more suitable for recommendation tasks. In contrast, the top-10 models in the LLM embedding benchmark MTEB  [^1] https://huggingface.co/spaces/mteb/leaderboard (Muennighoff et al., 2022), pretrained for various NLP tasks, use an average large embedding size of 4,506.
In this paper, we propose LLMInit, a novel initialization framework to bridge the embedding gaps between LLMs and CF models for effective and scalable recommendations. Specifically, we propose three strategies for selecting and transferring essential knowledge from LLM embeddings to CF models. For item embedding initialization, we explore the random selection, uniform selection, and variance-based selection to sample embedding index dimensions from LLM-generated representations. On the user side, where contextual information is frequently missing, we aggregate embeddings via a pooling operation over item neighborhoods. These approaches enable small CF models to selectively inherit the rich semantics of large language model embeddings while maintaining efficiency and scalability during the training and inference stage. Extensive experiments demonstrate the efficiency of LLMInit when applied to several state-of-the-art (SOTA) CF models and compared to other LLM-based models. We summarize our contributions as follows:
  • Conceptually, we identify the embedding collapse issue in the CF-based recommendations and propose a new paradigm of effectively leveraging LLMs for CF embedding initialization.
  • Methodologically, we devise three types of selective initialization strategies to inherit the rich world knowledge from pretrained LLMs into lightweight CF models for scalable recommendations.
  • Empirically, we implement LLMInit in a plug-and-play manner on various SOTA CF models, achieving significant performance and efficiency gains, especially compared to LLM-based RecSys.
2. Preliminaries
2.1. Problem Definition
Recommender systems (RecSys) aims to improve the platform’s accuracy in personalization by leveraging content features and user-item interactions. Formally, consider a textual bipartite graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ that models user-item interactions with node-level content features. The node set $\mathcal{V}$ is divided into users $u_{i}\in\mathcal{V}_{u}$ and items $v_{j}\in\mathcal{V}_{v}$. Edges $(u_{i},v_{j})\in\mathcal{E}$ represent interactions or relationships between users and items. The objective is to develop a recommendation algorithm that utilizes the node features and graph structure to predict and rank items that a user is likely to be interested in but has not yet interacted with.
Figure 2. An Illustration of LLMInit framework including contextual LLM input, three types of selective item embedding initialization, and the user embedding aggregation.
2.2. Related Work
LLMs for Recommendation. Recent advancements in this domain can be categorized into three paradigms (Wu et al., 2024): extracting embeddings for sequential recommendation (Qiu et al., 2021; Hou et al., 2022), generating semantic tokens to capture user preferences (Li et al., 2023; Xi et al., 2024), and directly adapting LLMs as RecSys through prompting (Hou et al., 2024; Liang et al., 2025) or tuning (Geng et al., 2022). However, they all fail to capture the intricate CF relationship in the user-item bipartite graphs as they either only focus on sequential order patterns (Qiu et al., 2021; Hou et al., 2022) or semantic understandings (Li et al., 2023; Xi et al., 2024; Hou et al., 2024; Zhang et al., 2023; Geng et al., 2022). In addition, they face notable challenges in efficiency and scalability due to the computational demands of LLMs, large-scale embeddings, overloaded vocabulary tokens, and limited in-context inputs in real-world scenarios. In contrast, our work LLMInit, proposes a novel paradigm to leverage LLMs in a scalable manner and addresses ineffective interaction modeling by resorting to CF models.
Table 1. Performance comparison of different CF models on four benchmark datasets in terms of Recall@10 and NDCG@10.
Method Beauty Toys-Games Tools-Home Office-Products
Recall@10 NDCG@10 Recall@10 NDCG@10 Recall@10 NDCG@10 Recall@10 NDCG@10
LightGCN (He et al., 2020) 0.0910 0.0432 0.0775 0.0360 0.0574 0.0283 0.0745 0.0365
+LLMInit-Rand 0.0960 (+5.5%) 0.0467 (+8.1%) 0.0805 (+3.9%) 0.0387 (+7.50%) 0.0612 (+6.6%) 0.0313 (+10.6%) 0.0773 (+3.8%) 0.0387 (+6.0%)
+LLMInit-Uni 0.1006 (+10.6%) 0.0469 (+8.6%) 0.0806 (+4.0%) 0.0388 (+7.8%) 0.0633 (+10.3%) 0.0319 (+12.7%) 0.0791 (+6.2%) 0.0395 (+8.2%)
+LLMInit-Var 0.1019 (+12.0%) 0.0485 (+12.3%) 0.0808 (+4.3%) 0.0389 (+8.1%) 0.0633 (+10.3%) 0.0317 (+12.0%) 0.0816 (+9.5%) 0.0414 (+13.4%)
SGL (Wu et al., 2021) 0.1017 0.0474 0.0832 0.0380 0.0580 0.0284 0.0669 0.0297
+LLMInit-Rand 0.1069 (+5.1%) 0.0520 (+9.7%) 0.0885 (+6.4%) 0.0418 (+10.0%) 0.0692 (+19.3%) 0.0337 (+18.7%) 0.0810 (+21.1%) 0.0426 (+43.4%)
+LLMInit-Uni 0.1101 (+8.3%) 0.0513 (+8.2%) 0.0920 (+10.6%) 0.0424 (+11.6%) 0.0676 (+16.6%) 0.0333 (+17.3%) 0.0773 (+15.6%) 0.0350 (+17.9%)
+LLMInit-Var 0.1106 (+8.8%) 0.0530 (+11.8%) 0.0927 (+11.4%) 0.0427 (+12.4%) 0.0686 (+18.3%) 0.0339 (+19.4%) 0.0794 (+18.7%) 0.0421 (+41.8%)
SGCL (Zhang et al., 2024b) 0.1027 0.0499 0.0828 0.0382 0.0585 0.0294 0.0647 0.0298
+LLMInit-Rand 0.1094 (+6.5%) 0.0512 (+2.6%) 0.0929 (+12.2%) 0.0418 (+9.4%) 0.0651 (+11.3%) 0.0326 (+10.9%) 0.0770 (+19.0%) 0.0365 (+22.5%)
+LLMInit-Uni 0.1115 (+8.6%) 0.0513 (+2.8%) 0.0923 (+11.5%) 0.0422 (+10.5%) 0.0650 (+11.1%) 0.0327 (+11.2%) 0.0742 (+14.7%) 0.0354 (+18.8%)
+LLMInit-Var 0.1104 (+7.5%) 0.0522 (+4.6%) 0.0941 (+13.6%) 0.0421 (+10.2%) 0.0646 (+10.4%) 0.0327 (+11.2%) 0.0776 (+19.9%) 0.0366 (+22.8%)
3. Method
As in Figure 2 LLMInit, before CF model training, the raw metadata is concatenated and fed to the LLMs. Then the semantic latent embeddings are selectively sampled via one of the CF item embedding initialization approaches followed by user embedding aggregation.
3.1. Selective Item Embedding Initialization
The CF item representation space can be considered a homomorphic projection of the sophisticated world-knowledge representation space in large language models (Sheng et al., 2024). Motivated by this, we propose to distill and selectively utilize the embeddings from LLMs to initialize the embeddings for CF recommendation models.
Formally, $\mathbf{v}\in\mathbb{R}^{N}$ represents the item embedding vector generated by the pre-trained LLMs. Let $K$ denotes the desired embedding dimensionality after selection. Then $\mathcal{I}\subseteq\{0,1,\dots,N-1\}$ is the selected indices. The resulting $K$-dimensional embedding is: $\mathbf{v}_{\text{selected}}=\{\mathbf{v}[i]:i\in\mathcal{I}\}.$ Here, we propose three novel initialization strategies to mitigate the inefficiency and embedding collapse issue for directly adopting the LLMs embeddings as follows.
3.1.1. Uniform Selection (LLMInit-Uni)
The uniform space among selections provides a balanced representation and helps mitigate potential redundancy or over-reliance on the LLM embeddings. By evenly spacing the indices, the strategy ensures that no region of the embedding vector dominates or is neglected across $\mathbf{v}$:
=$\mathcal{I}=\{k\cdot\lfloor N/K\rfloor:k\in\{0,1,\dots,K-1\}\}.$
3.1.2. Random Selection (LLMInit-Rand)
Randomization introduces diversity in the selected dimension patterns. Indices $K$ are randomly sampled from $\{0,1,\dots,N-1\}$ based on the uniform distribution:
$\mathcal{I}\sim\text{Random}(\{0,1,\dots,N-1\}),\quad|\mathcal{I}|=K.$
3.1.3. Variance Selection (LLMInit-Var)
Variance can serve as a heuristic for identifying information-rich and discriminative dimensions in the representation space. By selecting the top-$k$ dimensions of $\mathbf{v}$ with the highest variance across the dataset, this method prioritizes the most distinctive features, improving the separation of potential positive items from candidate pools. Let $\sigma^{2}[j]$ denote the variance of the $j$-th dimension of $\mathbf{h}$ across the dataset. The variance selects the embedding indices as:
$\mathcal{I}=\text{Top-}K(\sigma^{2}[0],\sigma^{2}[1],\dots,\sigma^{2}[N-1]).$
3.2. Aggregated User Embedding Initialization
Based on the item-side embedding, we design an aggregation strategy for user embedding initialization to tackle the privacy situation where the user’s context information is missing. Suppose we observe a user with historical items initial embeddings $\mathbf{v}_{j}$, and we design the aggregation initialization based on the smoothed neighborhood pooling process. The user embedding $\mathbf{u}_{i}$ is finalized as:
(1) $\mathbf{u}_{i}=\sum_{j\in N_{i}}\frac{1}{{|N_{i}|}}\mathbf{v}_{j},$
The normalization employs the degree ${|N_{i}|}$ to temper the magnitude and bias towards popular users after aggregation pooling.
4. Experiments
4.1. Experimental Set Up
Table 2. The statistics of the datasets.
Dataset # Users # Items # Interactions Sparsity
Beauty 10,554 6,087 94,148 99.8534%
Toys-Games 11,269 7,310 95,420 99.8842%
Tools-Home 9,246 6,199 73,250 99.8722%
Office-Products 3,261 1,584 36,544 99.2925%
We follow the 5-core setting in (Zhang et al., 2024b) and experiment on four Amazon  [^2] https://jmcauley.ucsd.edu/data/amazon/links.html real-world datasets: Beauty, Toys-Games, Tools-Home, and Office-Products, as summarized in Table  2. We use a leave-one-out strategy for data splitting and NDCG@10 and Recall@10 metrics for evaluation. For the CF baselines, we select LightGCN (He et al., 2020) as a representative of graph models, SGL (Wu et al., 2021) for self-supervised contrastive learning approaches, and SGCL (Zhang et al., 2024b) for the supervised contrastive learning paradigm. For LLM-based baselines, we use MoRec (Yuan et al., 2023) , LLMRank (Hou et al., 2024), and LLMRec (Wei et al., 2024). For fair comparisons, we set embedding size $K=128$ for all methods and LLM as MPNet (Song et al., 2020) for embedding generation unless specified.
4.2. Performance Comparison
The experimental results demonstrate that the proposed LLMInit strategies significantly enhance performance across all datasets compared to traditional baselines. Among all LLMInit strategies, LLMInit-Var consistently achieves the best results, leveraging variance-based initialization to capture CF and semantic nuances effectively. While LLMInit-Rand and LLMInit-Uni also outperform the baselines, the variance-based strategy emerges as the most robust and adaptable approach. Notably, in the Office-Products dataset, LLMInit delivers an average over 20% boost in NDCG@10 over SGL and SGCL. These findings underscore the robustness and effectiveness of LLMInit strategies, particularly LLMInit-Var, in enhancing recommendation and handling diverse datasets.
4.3. Efficiency Comparison of LLM-based RecSys
Table 3. Comparisons of computation cost and performance with LLM-based RecSys in full-ranking settings.
Methods Beauty Toys-Games
R@10 N@10 # Para. R@10 N@10 # Para.
MoRec (Yuan et al., 2023) 0.0857 0.0397 13M 0.0793 0.0362 14M
LLMRank (Hou et al., 2024) 7B 7B
LLMRec (Wei et al., 2024) 0.0974 0.0472 7B+13M 0.082 0.0389 7B+14M
LLMInit 0.1104 0.0522 0.2M 0.0941 0.0421 0.2M
Table 3 compares various LLM-based RecSys in full-ranking settings, focusing on the computation cost of parameter size and performances on the Beauty and Toys datasets. For the LLMRank (Hou et al., 2024) and LLMRec (Wei et al., 2024), we adopt the LLaMa-7B  [^3] https://huggingface.co/meta-llama/Llama-2-7b model for text generation. MoRec(Yuan et al., 2023), a modality-based model, leverages LLMs to generate textual embeddings. However, its performance is hindered by the use of a unified transform layer for embedding mapping to each item. Additionally, the parameter size for LLM-generated embeddings, initially around 13M/14M, can scale to billions as the total number of users and items surpasses 10 million, posing significant computational and storage challenges. Note that, LLMRank (Hou et al., 2024) directly employs an LLM for prompting but struggles with an incomplete long-context candidate pool and hallucination problem, leading to unreported performance results. LLMRec incorporates LLM for graph augmentation, improving recall to 0.0974 (Beauty) and 0.082 (Toys) but at a much higher computational cost (7B+ parameters). LLMInit, a lightweight alternative, achieves the best performance with only 0.2M parameters, highlighting the potential of parameter-efficient strategies in LLM-powered RecSys.
4.4. Cold-Start Recommendation
Figure 3. Performance improvement comparison in non-strict cold-start scenarios on Amazon-Beauty, where half of the observed interactions for each user are removed.
In the long-tail cold-start setting (Zhang et al., 2025), we remove half of the observed interactions of each user in training, resulting in many cold-start users with only a single interaction. Figure 3 highlights how the three CF models—LightGCN, SGL, and SGCL— react to the limited interactions and benefit from LLM-based initialization strategies in cold-start scenarios. LightGCN, as a simple graph model, is most resistant to the cold-start challenge, while the SGL and SGCL both suffer sharp performance drops due to reliance on interactions for data augmentation or supervised loss optimization. With the LLMInit, SGL demonstrates a stronger ability to utilize enhanced embeddings, by extending LightGCN with self-supervised loss. SGCL, which introduces supervised contrastive learning loss, achieves the highest gains with LLMInit strategies, with Recall and NDCG improving by over 50.4% and 58.9%, respectively. The supervised loss in SGCL amplifies the impact of informative embeddings, allowing it to extract more meaningful patterns from the selected dimensions. Overall, the results reveal that more advanced models like SGCL with nuanced supervised graph contrastive loss are better equipped to exploit high-quality LLM-based initializations.
4.5. The impacts of Large Language Models
Table 4. The Performance of LightGCN on Amazon-Beauty with initialization of different LLMs.
LLMs MPNet4441 gte-Qwen5552 GPT-S6663 GPT-L7773 Stella8884
Embedding Size 768 1,536 1,536 3,072 8,096
Rand R@10 0.0960 0.0936 0.0967 0.0925 0.0996
N@10 0.0467 0.0418 0.0463 0.044 0.0447
Uni R@10 0.1006 0.096 0.098 0.0942 0.0949
N@10 0.0469 0.0428 0.0465 0.0445 0.0417
Var R@10 0.1019 0.0932 0.098 0.096 0.0935
N@10 0.0485 0.0429 0.0465 0.046 0.0412
Table 4 illustrates the performance of LightGCN with different LLMs and initialization strategies (Rand, Uni, Var) on Amazon-Beauty. It observed that all LLMs can enhance LightGCN significantly with proper initialization strategies. Variance-based initialization (Var) generally outperforms random (Rand) and uniform (Uni) strategies by prioritizing high-variance dimensions, capturing the most informative features. Notably, MPNet, with the smallest embedding size (768), achieves the highest Recall@10 (0.1019) and NDCG@10 (0.0485), highlighting the importance of embedding quality over size. While larger models like GPT-L and Stella improve LightGCN with LLMInit, they still lag behind MPNet and GPT-S, emphasizing that LLM alignment with the dataset and strategic initialization matter more than embedding size alone.
5. Conclusion
In this work, we propose LLMInit, an effective and scalable framework that leverages pre-trained large language model embeddings for collaborative filtering model initialization. By employing selective strategies for item embeddings and aggregation pooling for user embeddings, LLMInit addresses data sparsity and cold-start challenges while ensuring scalability and efficiency with low-resource embedding size. Extensive experiments show LLMIint consistently enhancing diverse SOTA CF models, demonstrating LLMs’ new potential research direction to advance recommender systems.

### v2

@@
 LLMInit: A Free Lunch from Large Language Models for Selective Initialization of Recommendation
-Abstract.
-Collaborative filtering models, particularly graph-based approaches, have demonstrated strong performance in capturing user-item interactions for recommendation systems.
-However, they continue to struggle in cold-start and data-sparse scenarios.
-The emergence of large language models (LLMs) like GPT and LLaMA presents new possibilities for enhancing recommendation performance, especially in cold-start settings.
-Despite their promise, LLMs pose challenges related to scalability and efficiency due to their high computational demands and limited ability to model complex user-item relationships effectively.
-In this work, we introduce a novel perspective on leveraging LLMs for CF model initialization.
-Through experiments, we uncover an embedding collapse issue when scaling CF models to larger embedding dimensions.
-To effectively harness large-scale LLM embeddings, we propose innovative selective initialization strategies utilizing random, uniform, and variance-based index sampling.
-Our comprehensive evaluation on multiple real-world datasets demonstrates significant performance gains across various CF models while maintaining a lower computational cost compared to existing LLM-based recommendation approaches.
-Collaborative Filtering, Recommendation, Large Language Models
-doi: XXXXXXX.XXXXXXXccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
-1. Introduction
-Recommender systems (RecSys), particularly collaborative filtering (CF), play a pivotal role in many online platforms and web applications  (Koren et al., 2021; Fan et al., 2019; Zou et al., 2024), aiding users in navigating vast information by offering personalized suggestions.
-Among the various techniques, collaborative filtering (Wang et al., 2019; He et al., 2020; Xu et al., 2023; Zhang et al., 2024a; Xu et al., 2025) stands out for its effectiveness in predicting user preferences by analyzing past user-item interactions.
-A key principle among these models is learning randomly initialized user/item representations with carefully designed loss functions to capture each user’s preference (Rendle et al., 2012).
+Abstract
+Collaborative filtering (CF) is widely adopted in industrial recommender systems (RecSys) for modeling user-item interactions across numerous applications, but often struggles with cold-start and data-sparse scenarios.
+Recent advancements in pre-trained large language models (LLMs) with rich semantic knowledge, offer promising solutions to these challenges.
+However, deploying LLMs at scale is hindered by their significant computational demands and latency.
+In this paper, we propose a novel and scalable LLM-RecSys framework, LLMInit, designed to integrate pretrained LLM embeddings into CF models through selective initialization strategies.
+Specifically, we identify the embedding collapse issue observed when CF models scale and match the large embedding sizes in LLMs and avoid the problem by introducing efficient sampling methods, including, random, uniform, and variance-based selections.
+Comprehensive experiments conducted on multiple real-world datasets demonstrate that LLMInit significantly improves recommendation performance while maintaining low computational costs, offering a practical and scalable solution for industrial applications.
+To facilitate industry adoption and promote future research, we provide open-source access to our implementation at https://github.com/DavidZWZ/LLMInit.
+LLMInit: A Free Lunch from Large Language Models
+for Selective Initialization of Recommendation
+Weizhi Zhang1, Liangwei Yang2thanks: Corresponding author: liangwei.yang@salesforce.com, Wooseong Yang1, Henry Peng Zou1, Yuqing Liu1, Ke Xu1, Sourav Medya 1, Philip S.
+Yu1, 1University of Illinois Chicago, 2Salesforce AI Research
+1 Introduction
+Figure 1: An investigation of the embedding collapse issue of CF model LightGCN he2020lightgcn in two Amazon-review datasets.
+Increasing embedding size with exponentially growing parameters finally leads to performance degradation.
+Recommender systems (RecSys), particularly collaborative filtering (CF), play a pivotal role in many online platforms and web applications  koren2021advances; fan2019graph, aiding users in navigating vast information by offering personalized suggestions.
+Among the various techniques, collaborative filtering wang2019neural; he2020lightgcn; zhang2024we stands out for its effectiveness in predicting user preferences by analyzing past user-item interactions.
+A key principle among these models is learning randomly initialized user/item representations with carefully designed loss functions to capture each user’s preference rendle2012bpr.
 However, such CF approaches heavily rely on the observed user-item interactions to optimize the CF embeddings from random initialization and perform poorly in real-world applications where interaction data is sparse and users/items are cold-start.
-More recently, the rapid advancement of large language models (LLMs) such as GPT-4  (Achiam et al., 2023) and LLaMa  (Touvron et al., 2023), which have exhibited remarkable proficiency in understanding and processing textual information, have piqued the interest of researchers seeking to explore their potential in improving recommender systems beyond traditional approaches  (Geng et al., 2022; Yuan et al., 2023; Li et al., 2023).
+More recently, the rapid advancement of large language models (LLMs) such as GPT-4  achiam2023gpt and LLaMa  touvron2023llama, which have exhibited remarkable proficiency in understanding and processing textual information, have piqued the interest of researchers seeking to explore their potential in improving recommender systems beyond traditional approaches  geng2022recommendation; li2023text; zhang2025personaagent.
 One of the most promising directions in this exploration involves adapting LLMs as recommender systems through prompt engineering and tailored instruction tuning.
-Such integration of LLMs enables the RecSys to better understand complex user preferences, and interpret contextual information, thereby surpassing the traditional models in cold-start settings.
-Figure 1.
-An investigation of the embedding collapse issue of CF model LightGCN (He et al., 2020) in two Amazon datasets.
-Increasing embedding size with exponentially growing parameters finally leads to performance degradation.
+Such integration of LLMs enables the RecSys to better understand user preferences, and interpret contextual information, thereby surpassing the traditional models in cold-start settings.
 Despite the inspiring progress made in LLMs for recommendations, their adoption in online deployments faces significant challenges related to efficiency and scalability.
@@
 Such problems can be further exaggerated in real-world applications of large-scale users and items due to the enlarged vocabulary tokens and in-context input.
-Moreover, LLMs fall short of complex CF interaction understanding owing to the next-token prediction pipeline and the constraints of a maximum token limit (e.g., 4096 tokens in LLaMA) to incorporate the whole CF interactions.
-This limitation hampers their ability to capture and effectively model global user-item dependencies as in CF models, leading to poor performance in the full-ranking (Hou et al., 2024) and warm-start settings (Bao et al., 2023).
+Moreover, LLMs fall short of complex CF interaction understanding owing to the next-token prediction pipeline and the constraints of a maximum token limit (e.g., 4096 tokens in LLaMA-2) to incorporate the whole CF interactions.
+This limitation hampers their ability to capture and effectively model global user-item dependencies as in CF models, leading to poor performance in the full-ranking hou2024large and warm-start settings bao2023tallrec.
 Therefore, we raise a critical research question:
-How can we harness the power of LLMs for recommendation systems in an effective and scalable manner?
+How can we harness the power of LLMs for RecSys in an effective and scalable manner?
 To answer the above question, we propose to utilize LLMs for CF-based RecSys from a new perspective by taking pretrained large language models as a free lunch for the embedding initialization of CF models.
@@
 However, storing the large-scale embeddings from LLMs for each user and item is not efficient nor scalable, especially when the platform scales to millions of users and items.
-Furthermore, as illustrated in Figure 1, unlike the scaling laws observed in NLP models (Kaplan et al., 2020), the performance of the recommendation model LightGCN (He et al., 2020) deteriorates even when the numbers of parameters scale up exponentially.
-This behavior resembles the embedding collapse issue (Guo et al., [n.
-d.]) observed in click-through rate prediction models (Guo et al., 2017; Wang et al., 2021).
+Furthermore, as illustrated in Figure 1, unlike the scaling laws observed in language models kaplan2020scaling and graph models liu2024towards, the performance of the recommendation model LightGCN he2020lightgcn deteriorates even when the numbers of parameters scale up exponentially.
+This behavior resembles the embedding collapse issue guo2024embedding observed in click-through rate prediction models wang2021dcn.
 These findings suggest that significantly smaller embedding sizes (e.g., 128) are more suitable for recommendation tasks.
-In contrast, the top-10 models in the LLM embedding benchmark MTEB  [^1] https://huggingface.co/spaces/mteb/leaderboard (Muennighoff et al., 2022), pretrained for various NLP tasks, use an average large embedding size of 4,506.
-In this paper, we propose LLMInit, a novel initialization framework to bridge the embedding gaps between LLMs and CF models for effective and scalable recommendations.
-Specifically, we propose three strategies for selecting and transferring essential knowledge from LLM embeddings to CF models.
-For item embedding initialization, we explore the random selection, uniform selection, and variance-based selection to sample embedding index dimensions from LLM-generated representations.
-On the user side, where contextual information is frequently missing, we aggregate embeddings via a pooling operation over item neighborhoods.
-These approaches enable small CF models to selectively inherit the rich semantics of large language model embeddings while maintaining efficiency and scalability during the training and inference stage.
-Extensive experiments demonstrate the efficiency of LLMInit when applied to several state-of-the-art (SOTA) CF models and compared to other LLM-based models.
-We summarize our contributions as follows:
+In contrast, the top-10 models in the LLM embedding benchmark MTEB muennighoff2023mteb, pretrained for various NLP tasks, use an average large embedding size of 4,506.
+Figure 2: An Illustration of LLMInit framework including contextual LLM input, three types of selective item embedding initialization strategies, and the user embedding aggregation operation.
+In this paper, we present LLMInit, a practical and scalable initialization framework that bridges large language model (LLM) embeddings with collaborative filtering (CF) models for industrial-scale recommendation systems.
+LLMInit tackles the deployment challenges of LLMs and enhances the semantic capacity of CF models by initializing CF embedding models based on pretrained LLMs.
+To this end, we introduce three efficient strategies, including random, uniform, and variance-based index selection, that selectively sample and compress LLM-generated item embeddings to suit the constrained embedding spaces of CF models.
+On the user side, where contextual signals may be missing or sparse, LLMInit aggregates item-level embeddings using lightweight pooling mechanisms.
+This design enables recommendation models to inherit rich language semantics from pretrained LLMs while maintaining high efficiency in both training and inference.
+Extensive evaluations on multiple real-world datasets demonstrate that LLMInit consistently enhances performance across several state-of-the-art CF architectures.
+Compared to full LLM-based recommendation systems, our framework achieves significant gains in scalability and efficiency, making it well-suited for real-world deployments.
+Our contributions underscore the value of adapting language model innovations to meet the practical demands of large-scale, industry-grade recommendation systems.
   • Conceptually, we identify the embedding collapse issue in the CF-based recommendations and propose a new paradigm of effectively leveraging LLMs for CF embedding initialization.
-  • Methodologically, we devise three types of selective initialization strategies to inherit the rich world knowledge from pretrained LLMs into lightweight CF models for scalable recommendations.
-  • Empirically, we implement LLMInit in a plug-and-play manner on various SOTA CF models, achieving significant performance and efficiency gains, especially compared to LLM-based RecSys.
-2. Preliminaries
-2.1. Problem Definition
+  • Methodologically, we devise three types of selective initialization strategies to inherit the rich pretrained knowledge from LLMs into lightweight CF models for scalable recommendations.
+  • Empirically, we implement LLMInit in a plug-and-play manner on various SOTA CF models, achieving significant performance and efficiency gains, especially compared to LLM RecSys.
+2 Preliminaries
+2.1 Problem Definition
 Recommender systems (RecSys) aims to improve the platform’s accuracy in personalization by leveraging content features and user-item interactions.
@@
 The objective is to develop a recommendation algorithm that utilizes the node features and graph structure to predict and rank items that a user is likely to be interested in but has not yet interacted with.
-Figure 2. An Illustration of LLMInit framework including contextual LLM input, three types of selective item embedding initialization, and the user embedding aggregation.
-2.2. Related Work
+2.2 Related Work
 LLMs for Recommendation.
-Recent advancements in this domain can be categorized into three paradigms (Wu et al., 2024): extracting embeddings for sequential recommendation (Qiu et al., 2021; Hou et al., 2022), generating semantic tokens to capture user preferences (Li et al., 2023; Xi et al., 2024), and directly adapting LLMs as RecSys through prompting (Hou et al., 2024; Liang et al., 2025) or tuning (Geng et al., 2022).
-However, they all fail to capture the intricate CF relationship in the user-item bipartite graphs as they either only focus on sequential order patterns (Qiu et al., 2021; Hou et al., 2022) or semantic understandings (Li et al., 2023; Xi et al., 2024; Hou et al., 2024; Zhang et al., 2023; Geng et al., 2022).
+Recent advancements in this domain can be categorized into three paradigms wu2024survey: extracting embeddings for sequential recommendation qiu2021u; hou2022towards, generating semantic tokens to capture user preferences li2023text; xi2024towards, and directly adapting LLMs as RecSys through prompting hou2024large; liang-etal-2025-taxonomy or tuning geng2022recommendation; yang2024item.
+However, they all fail to capture the intricate CF relationship in the user-item bipartite graphs as they either only focus on sequential order patterns qiu2021u; hou2022towards or semantic understandings li2023text; xi2024towards; zhang2023dual.
 In addition, they face notable challenges in efficiency and scalability due to the computational demands of LLMs, large-scale embeddings, overloaded vocabulary tokens, and limited in-context inputs in real-world scenarios.
 In contrast, our work LLMInit, proposes a novel paradigm to leverage LLMs in a scalable manner and addresses ineffective interaction modeling by resorting to CF models.
-Table 1. Performance comparison of different CF models on four benchmark datasets in terms of Recall@10 and NDCG@10.
-Method Beauty Toys-Games Tools-Home Office-Products
-Recall@10 NDCG@10 Recall@10 NDCG@10 Recall@10 NDCG@10 Recall@10 NDCG@10
-LightGCN (He et al., 2020) 0.0910 0.0432 0.0775 0.0360 0.0574 0.0283 0.0745 0.0365
-+LLMInit-Rand 0.0960 (+5.5%) 0.0467 (+8.1%) 0.0805 (+3.9%) 0.0387 (+7.50%) 0.0612 (+6.6%) 0.0313 (+10.6%) 0.0773 (+3.8%) 0.0387 (+6.0%)
-+LLMInit-Uni 0.1006 (+10.6%) 0.0469 (+8.6%) 0.0806 (+4.0%) 0.0388 (+7.8%) 0.0633 (+10.3%) 0.0319 (+12.7%) 0.0791 (+6.2%) 0.0395 (+8.2%)
-+LLMInit-Var 0.1019 (+12.0%) 0.0485 (+12.3%) 0.0808 (+4.3%) 0.0389 (+8.1%) 0.0633 (+10.3%) 0.0317 (+12.0%) 0.0816 (+9.5%) 0.0414 (+13.4%)
-SGL (Wu et al., 2021) 0.1017 0.0474 0.0832 0.0380 0.0580 0.0284 0.0669 0.0297
-+LLMInit-Rand 0.1069 (+5.1%) 0.0520 (+9.7%) 0.0885 (+6.4%) 0.0418 (+10.0%) 0.0692 (+19.3%) 0.0337 (+18.7%) 0.0810 (+21.1%) 0.0426 (+43.4%)
-+LLMInit-Uni 0.1101 (+8.3%) 0.0513 (+8.2%) 0.0920 (+10.6%) 0.0424 (+11.6%) 0.0676 (+16.6%) 0.0333 (+17.3%) 0.0773 (+15.6%) 0.0350 (+17.9%)
-+LLMInit-Var 0.1106 (+8.8%) 0.0530 (+11.8%) 0.0927 (+11.4%) 0.0427 (+12.4%) 0.0686 (+18.3%) 0.0339 (+19.4%) 0.0794 (+18.7%) 0.0421 (+41.8%)
-SGCL (Zhang et al., 2024b) 0.1027 0.0499 0.0828 0.0382 0.0585 0.0294 0.0647 0.0298
-+LLMInit-Rand 0.1094 (+6.5%) 0.0512 (+2.6%) 0.0929 (+12.2%) 0.0418 (+9.4%) 0.0651 (+11.3%) 0.0326 (+10.9%) 0.0770 (+19.0%) 0.0365 (+22.5%)
-+LLMInit-Uni 0.1115 (+8.6%) 0.0513 (+2.8%) 0.0923 (+11.5%) 0.0422 (+10.5%) 0.0650 (+11.1%) 0.0327 (+11.2%) 0.0742 (+14.7%) 0.0354 (+18.8%)
-+LLMInit-Var 0.1104 (+7.5%) 0.0522 (+4.6%) 0.0941 (+13.6%) 0.0421 (+10.2%) 0.0646 (+10.4%) 0.0327 (+11.2%) 0.0776 (+19.9%) 0.0366 (+22.8%)
-3. Method
+Table 1: Performance comparison on four datasets. Gains are relative to the base model’s performance.
+[…]
+3 Method
 As in Figure 2 LLMInit, before CF model training, the raw metadata is concatenated and fed to the LLMs.
 Then the semantic latent embeddings are selectively sampled via one of the CF item embedding initialization approaches followed by user embedding aggregation.
-3.1. Selective Item Embedding Initialization
-The CF item representation space can be considered a homomorphic projection of the sophisticated world-knowledge representation space in large language models (Sheng et al., 2024).
+3.1 Selective Item Embedding Initialization
+The CF item representation space can be regarded as a subspace of the sophisticated world-knowledge representation space in large language models sheng2024language.
 Motivated by this, we propose to distill and selectively utilize the embeddings from LLMs to initialize the embeddings for CF recommendation models.
-Formally, $\mathbf{v}\in\mathbb{R}^{N}$ represents the item embedding vector generated by the pre-trained LLMs.
+Formally, $\mathbf{v}\in\mathbf{R}^{N}$ represents the item embedding vector generated by the pre-trained LLMs.
 Let $K$ denotes the desired embedding dimensionality after selection.
@@
 The resulting $K$-dimensional embedding is: $\mathbf{v}_{\text{selected}}=\{\mathbf{v}[i]:i\in\mathcal{I}\}.$ Here, we propose three novel initialization strategies to mitigate the inefficiency and embedding collapse issue for directly adopting the LLMs embeddings as follows.
-3.1.1. Uniform Selection (LLMInit-Uni)
+3.1.1 Uniform Selection (LLMInit-Uni)
 The uniform space among selections provides a balanced representation and helps mitigate potential redundancy or over-reliance on the LLM embeddings.
 By evenly spacing the indices, the strategy ensures that no region of the embedding vector dominates or is neglected across $\mathbf{v}$:
-=$\mathcal{I}=\{k\cdot\lfloor N/K\rfloor:k\in\{0,1,\dots,K-1\}\}.$
-3.1.2. Random Selection (LLMInit-Rand)
+$\mathcal{I}=\{k\cdot\lfloor N/K\rfloor:k\in\{0,1,\dots,K-1\}\}.$
+3.1.2 Random Selection (LLMInit-Rand)
 Randomization introduces diversity in the selected dimension patterns. Indices $K$ are randomly sampled from $\{0,1,\dots,N-1\}$ based on the uniform distribution:
 $\mathcal{I}\sim\text{Random}(\{0,1,\dots,N-1\}),\quad|\mathcal{I}|=K.$
-3.1.3. Variance Selection (LLMInit-Var)
+3.1.3 Variance Selection (LLMInit-Var)
 Variance can serve as a heuristic for identifying information-rich and discriminative dimensions in the representation space.
@@
 $\mathcal{I}=\text{Top-}K(\sigma^{2}[0],\sigma^{2}[1],\dots,\sigma^{2}[N-1]).$
-3.2. Aggregated User Embedding Initialization
+3.2 Aggregated User Embedding Initialization
 Based on the item-side embedding, we design an aggregation strategy for user embedding initialization to tackle the privacy situation where the user’s context information is missing.
@@
 The user embedding $\mathbf{u}_{i}$ is finalized as:
-(1) $\mathbf{u}_{i}=\sum_{j\in N_{i}}\frac{1}{{|N_{i}|}}\mathbf{v}_{j},$
+$\mathbf{u}_{i}=\sum_{j\in N_{i}}\frac{1}{{|N_{i}|}}\mathbf{v}_{j},$ (1)
 The normalization employs the degree $|N_{i}|$} to temper the magnitude and bias towards popular users after aggregation pooling.
-4. Experiments
-4.1. Experimental Set Up
-Table 2. The statistics of the datasets.
-Dataset # Users # Items # Interactions Sparsity
-Beauty 10,554 6,087 94,148 99.8534%
-Toys-Games 11,269 7,310 95,420 99.8842%
-Tools-Home 9,246 6,199 73,250 99.8722%
-Office-Products 3,261 1,584 36,544 99.2925%
-We follow the 5-core setting in (Zhang et al., 2024b) and experiment on four Amazon  [^2] https://jmcauley.ucsd.edu/data/amazon/links.html real-world datasets: Beauty, Toys-Games, Tools-Home, and Office-Products, as summarized in Table  2.
-We use a leave-one-out strategy for data splitting and NDCG@10 and Recall@10 metrics for evaluation.
-For the CF baselines, we select LightGCN (He et al., 2020) as a representative of graph models, SGL (Wu et al., 2021) for self-supervised contrastive learning approaches, and SGCL (Zhang et al., 2024b) for the supervised contrastive learning paradigm.
-For LLM-based baselines, we use MoRec (Yuan et al., 2023) , LLMRank (Hou et al., 2024), and LLMRec (Wei et al., 2024).
-For fair comparisons, we set embedding size $K=128$ for all methods and LLM as MPNet (Song et al., 2020) for embedding generation unless specified.
-4.2. Performance Comparison
-The experimental results demonstrate that the proposed LLMInit strategies significantly enhance performance across all datasets compared to traditional baselines.
-Among all LLMInit strategies, LLMInit-Var consistently achieves the best results, leveraging variance-based initialization to capture CF and semantic nuances effectively.
-While LLMInit-Rand and LLMInit-Uni also outperform the baselines, the variance-based strategy emerges as the most robust and adaptable approach.
-Notably, in the Office-Products dataset, LLMInit delivers an average over 20% boost in NDCG@10 over SGL and SGCL.
-These findings underscore the robustness and effectiveness of LLMInit strategies, particularly LLMInit-Var, in enhancing recommendation and handling diverse datasets.
-4.3. Efficiency Comparison of LLM-based RecSys
-Table 3. Comparisons of computation cost and performance with LLM-based RecSys in full-ranking settings.
+4 Experiments
+4.1 Experimental Set Up
+We adopt the 5-core setting from zhang2025sgcl; zhang2024mixed and conduct experiments on four real-world Amazon datasets ni2019justifying,  [^1] https://jmcauley.ucsd.edu/data/amazon/links.html, including Beauty, Toys-Games, Tools-Home, and Office-Products, as summarized in Table 4.
+We use a leave-one-out strategy for data splitting and evaluate all the models using top-K metrics for ranking, including Recall@10 and NDCG@10.
+For collaborative filtering (CF) baselines, we include LightGCN he2020lightgcn as a representative graph-based model, SGL wu2021self for self-supervised contrastive learning, and SGCL zhang2025sgcl for supervised contrastive learning.
+For LLM-based baselines, we evaluate MoRec yuan2023go, LLMRank hou2024large, LLMRec wei2024llmrec, and TIGER rajput2023recommender.
+To ensure fair comparison, we fix the embedding dimension to $K=128$ across all models and use MPNet song2020mpnet as the default embedding generator.
+Table 2: Comparisons of computation cost and performance with LLM-based RecSys in full-ranking settings.
 Methods Beauty Toys-Games
 R@10 N@10 # Para. R@10 N@10 # Para.
-MoRec (Yuan et al., 2023) 0.0857 0.0397 13M 0.0793 0.0362 14M
-LLMRank (Hou et al., 2024) 7B 7B
-LLMRec (Wei et al., 2024) 0.0974 0.0472 7B+13M 0.082 0.0389 7B+14M
-LLMInit 0.1104 0.0522 0.2M 0.0941 0.0421 0.2M
-Table 3 compares various LLM-based RecSys in full-ranking settings, focusing on the computation cost of parameter size and performances on the Beauty and Toys datasets.
-For the LLMRank (Hou et al., 2024) and LLMRec (Wei et al., 2024), we adopt the LLaMa-7B  [^3] https://huggingface.co/meta-llama/Llama-2-7b model for text generation.
-MoRec(Yuan et al., 2023), a modality-based model, leverages LLMs to generate textual embeddings.
-However, its performance is hindered by the use of a unified transform layer for embedding mapping to each item.
-Additionally, the parameter size for LLM-generated embeddings, initially around 13M/14M, can scale to billions as the total number of users and items surpasses 10 million, posing significant computational and storage challenges.
-Note that, LLMRank (Hou et al., 2024) directly employs an LLM for prompting but struggles with an incomplete long-context candidate pool and hallucination problem, leading to unreported performance results.
-LLMRec incorporates LLM for graph augmentation, improving recall to 0.0974 (Beauty) and 0.082 (Toys) but at a much higher computational cost (7B+ parameters).
-LLMInit, a lightweight alternative, achieves the best performance with only 0.2M parameters, highlighting the potential of parameter-efficient strategies in LLM-powered RecSys.
-4.4. Cold-Start Recommendation
-Figure 3. Performance improvement comparison in non-strict cold-start scenarios on Amazon-Beauty, where half of the observed interactions for each user are removed.
-In the long-tail cold-start setting (Zhang et al., 2025), we remove half of the observed interactions of each user in training, resulting in many cold-start users with only a single interaction.
-Figure 3 highlights how the three CF models—LightGCN, SGL, and SGCL— react to the limited interactions and benefit from LLM-based initialization strategies in cold-start scenarios.
+MoRec yuan2023go 0.0857 0.0397 13M 0.0793 0.0362 14M
+LLMRank hou2024large 7B 7B
+LLMRec wei2024llmrec 0.0974 0.0472 7B+13M 0.0820 0.0389 7B+14M
+TIGER rajput2023recommender 0.1015 0.0481 13M 0.0864 0.0385 13M
+LLMInit-Full 0.1029 0.0475 13M 0.0871 0.0387 14M
+LLMInit-Var 0.1104 0.0522 2M 0.0941 0.0421 2M
+4.2 Performance Comparison
+The experimental results clearly demonstrate that all variants of LLMInit substantially improve recommendation performance across diverse datasets compared to traditional CF baselines.
+Among the proposed strategies, LLMInit-Var consistently achieves the best results, effectively capturing both collaborative filtering signals and semantic information through variance-based index selection.
+While both LLMInit-Rand and LLMInit-Uni offer notable gains, the variance-based method proves to be the most robust and adaptable across model types and domains.
+Notably, on the Office-Products dataset (with severe embedding collapse issue as in Figuire 1), LLMInit-Var delivers more than a 20% average gain in NDCG@10 over strong baselines like SGL and SGCL.
+These consistent improvements highlight the generalizability and practicality of LLMInit, especially in handling diverse real-world scenarios.
+4.3 Efficiency and Perforamance Comparisons of LLM-based RecSys
+Table 2 compares computation cost and performance of various LLM-based RecSys in full-ranking settings.
+For the LLMRank hou2024large and LLMRec wei2024llmrec, we adopt the LLaMa-7B model for text generation.
+LLMRank hou2024large directly employs an LLM for prompting but struggles with an incomplete long-context candidate pool and hallucination problem, leading to unreported performance results.
+LLMRec prompts LLM for graph augmentation, achieving recall 0.0974 (Beauty) and 0.082 (Toys) with high computational cost (7B+ parameters).
+In comparison, MoRecyuan2023go, a modality-based model, leverages LLMs to generate textual embeddings but restricted by the use of a unified transform layer for distinct item embedding mapping.
+In addition, the parameter size for LLM-generated embeddings, initially around 13M/14M for MoRec and LLMInit-Full, can scale to billions as the total number of users and items surpasses 1 million, posing significant computational and storage challenges.
+TIGER, employing vector quantization for generative recommendation, shows potential for scalable RecSys.
+LLMInit-Var, a lightweight alternative, achieves the best performance with only 2M parameters, highlighting the potential of parameter-efficient strategies in LLM-based RecSys.
+Figure 3: Performance improvement comparison in non-strict cold-start scenarios on Amazon-Beauty, where half of the observed interactions for each user are removed.
+Table 3: The Performance of LightGCN on Amazon-Beauty with initialization of different LLMs.
+The embedding size of LLMs scale from 768 to 8,096 while LLMInit models all adopt the embedding size of 128, efficient for large user base scenarios.
+LLMs MPNet2222 gte-Qwen3333 GPT-S4444 GPT-L5555 Stella6666
+Embedding Size 768 1,536 1,536 3,072 8,096
+LLMInit-Rand Recall@10 0.0960 0.0936 0.0967 0.0925 0.0996
+NDCG@10 0.0467 0.0418 0.0463 0.044 0.0447
+LLMInit-Uni Recall@10 0.1006 0.096 0.098 0.0942 0.0949
+NDCG@10 0.0469 0.0428 0.0465 0.0445 0.0417
+LLMInit-Var Recall@10 0.1019 0.0932 0.098 0.096 0.0935
+NDCG@10 0.0485 0.0429 0.0465 0.046 0.0412
+4.4 Cold-Start Recommendation
+In the long-tail cold-start setting zhang2025cold; yang2025cold, we remove half of the observed interactions of each user in training, resulting in many cold-start users with only a single interaction.
+Figure 3 highlights how the three CF models (LightGCN, SGL, and SGCL) react to the limited interactions and benefit from LLM-based initialization strategies in cold-start scenarios.
 LightGCN, as a simple graph model, is most resistant to the cold-start challenge, while the SGL and SGCL both suffer sharp performance drops due to reliance on interactions for data augmentation or supervised loss optimization.
-With the LLMInit, SGL demonstrates a stronger ability to utilize enhanced embeddings, by extending LightGCN with self-supervised loss.
-SGCL, which introduces supervised contrastive learning loss, achieves the highest gains with LLMInit strategies, with Recall and NDCG improving by over 50.4% and 58.9%, respectively.
+With the LLMInit, SGL and SGCL demonstrate stronger abilities to utilize enhanced embeddings.
+In particular, SGCL achieves the highest gains with LLMInit strategies, with Recall and NDCG improving by over 50.4% and 58.9%, respectively.
 The supervised loss in SGCL amplifies the impact of informative embeddings, allowing it to extract more meaningful patterns from the selected dimensions.
 Overall, the results reveal that more advanced models like SGCL with nuanced supervised graph contrastive loss are better equipped to exploit high-quality LLM-based initializations.
-4.5. The impacts of Large Language Models
-Table 4. The Performance of LightGCN on Amazon-Beauty with initialization of different LLMs.
-LLMs MPNet4441 gte-Qwen5552 GPT-S6663 GPT-L7773 Stella8884
-Embedding Size 768 1,536 1,536 3,072 8,096
-Rand R@10 0.0960 0.0936 0.0967 0.0925 0.0996
-N@10 0.0467 0.0418 0.0463 0.044 0.0447
-Uni R@10 0.1006 0.096 0.098 0.0942 0.0949
-N@10 0.0469 0.0428 0.0465 0.0445 0.0417
-Var R@10 0.1019 0.0932 0.098 0.096 0.0935
-N@10 0.0485 0.0429 0.0465 0.046 0.0412
-Table 4 illustrates the performance of LightGCN with different LLMs and initialization strategies (Rand, Uni, Var) on Amazon-Beauty.
-It observed that all LLMs can enhance LightGCN significantly with proper initialization strategies.
-Variance-based initialization (Var) generally outperforms random (Rand) and uniform (Uni) strategies by prioritizing high-variance dimensions, capturing the most informative features.
-Notably, MPNet, with the smallest embedding size (768), achieves the highest Recall@10 (0.1019) and NDCG@10 (0.0485), highlighting the importance of embedding quality over size.
-While larger models like GPT-L and Stella improve LightGCN with LLMInit, they still lag behind MPNet and GPT-S, emphasizing that LLM alignment with the dataset and strategic initialization matter more than embedding size alone.
-5. Conclusion
-In this work, we propose LLMInit, an effective and scalable framework that leverages pre-trained large language model embeddings for collaborative filtering model initialization.
-By employing selective strategies for item embeddings and aggregation pooling for user embeddings, LLMInit addresses data sparsity and cold-start challenges while ensuring scalability and efficiency with low-resource embedding size.
-Extensive experiments show LLMIint consistently enhancing diverse SOTA CF models, demonstrating LLMs’ new potential research direction to advance recommender systems.
+4.5 The impacts of Large Language Models
+Table 3 presents the performance of LightGCN initialized with embeddings from various LLMs using different strategies on the Amazon-Beauty dataset.
+The results show that all tested LLMs can enhance LightGCN’s performance when paired with appropriate initialization methods.
+Among them, the variance-based strategy consistently delivers superior results by selecting high-variance dimensions that retain the most informative semantic features.
+While larger models such as GPT-L and Stella provide higher-dimensional embeddings, their advantages are not linearly correlated with model size.
+Instead, MPNet and GPT-S exhibit superior performance, suggesting that the quality and domain alignment of embeddings play a more decisive role than sheer model capacity.
+Consequently, optimizing initialization and embedding relevance emerges as a more effective approach than simply increasing embedding dimensionality or model size for large-scale recommendation scenarios.
+5 Conclusion
+In this work, we introduced LLMInit, a practical and scalable framework designed for industrial LLM RecSys, which effectively integrates LLM embeddings for CF model initialization.
+By employing selective strategies for item embeddings and aggregation pooling for user embeddings, LLMInit addresses critical industry challenges of existing CF models, such as data sparsity and cold-start scenarios, while maintaining computational efficiency suitable for large-scale deployments.
+Extensive evaluations on real-world datasets validate that LLMInit significantly boosts the performance of various state-of-the-art CF models, demonstrating the tangible value and feasibility of leveraging language-based world knowledge to improve recommendation quality in industrial applications.
+Acknowledgment
+This work is supported in part by NSF under grants III-2106758, and POSE-2346158.
+Limitations
+Although LLMInit achieves notable improvements for scalable LLM-based recommendation, especially under cold-start and sparse-data scenarios, its effectiveness relies on the quality and domain relevance of pretrained LLM embeddings, necessitating careful selection of LLMs for domains or conducting further domain adaptation.
+Additionally, our current approach leverages only textual metadata; incorporating multi-modal features (such as images or audio) remains an important future direction to enhance real-world applicability.
+Finally, more advanced CF models such ordinary-differential-equations (ODEs) xu2023graph; xu2025graph; zhang2024we have not yet been evaluated within LLMInit.
+Appendix A Dataset Statistics
+Table 4: Statistics of the datasets.
+Dataset # Users # Items # Inter. Sparsity Beauty 10,554 6,087 94,148 99.85% Toys–Games 11,269 7,310 95,420 99.88% Tools–Home 9,246 6,199 73,250 99.87% Office–Products 3,261 1,584 36,544 99.29%
+To comprehensively evaluate the effectiveness of LLM-based initialization strategies in real-world recommendation scenarios, we conduct experiments on four widely used Amazon product review datasets (ni2019justifying), covering diverse user interaction domains.
+Each dataset consists of implicit user–item interactions, where an interaction denotes a user’s engagement (e.g., rating and review) with a product.
+Following common practice, we filter out users and items with fewer than five interactions and adopt the leave-one-out evaluation protocol.
+The statistics of the processed datasets are summarized in Table 4.
+Appendix B Pooling Strategies
+Table 5: Performance comparison of different initialization pooling strategies on Amazon-Beauty. All models are initialized with LLM-based embeddings except for the random baseline.
+Method Recall@10 NDCG@10
+Random Init 0.1029 0.0475
+Mean (LLMInit) 0.1104 0.0522
+Max (LLMInit) 0.1088 0.0510
+LightGCN (LLMInit) 0.1101 0.0515
+To further analyze the impact of different embedding aggregation strategies, Table 5 compares pooling-based methods for user embedding initializations.
+Results show that all LLM-based initialization strategies outperform random initialization, confirming the benefits of semantic priors learned by LLMs.
+Mean pooling achieves the best overall performance, slightly surpassing both max pooling and LightGCN pooling.
+This indicates that averaging item embeddings provides a stable and representative initialization signal for graph-based user–item interaction modeling.